1. How many particles are in 125 moles of Sodium Acetate (NaC2H3O2)
1
Expert's answer
2021-01-15T06:31:50-0500

A mole of Sodium Acetate contains 6.02×1023 molecules.

Proportion:

1 mol - 6.02×1023 molecules

125 mol - x

"x = \frac{125 \times 6.02 \times 10^{23}}{1}= 7.52 \times 10^{25} \;molecules"

Answer: 7.52×1025 molecules